The perimeter of a rectangle is 70 m. What are the dimensions that will produce the maximum area of such a rectangle?
Answer by checkley77(12844)

70/4=17.5m. for each side of a square should do it.
17.5*17.5=306.25 is the maximum area possible.
Proof:
Try 17.4 & 17.6
17.4*17.6=306.24
17*18=306